x/2 + y/4 = 1/7
x/4 + y/6 = 1/4
solve for x and y
I got x = 29/49 if this is correct please help me determine y.
thank you.

Answer by fractalier(6550)

From
x/2 + y/4 = 1/7
x/4 + y/6 = 1/4
we can multiply the bottom one by two and subtract...we get
x/2 + y/4 = 1/7
-(x/2 + y/3 = 1/2)
and then have
-y/12 = -5/14
so that
y = 30/7
now plug that in to either equation and get
x/2 + 15/14 = 1/7
multiply all by 14
7x + 15 = 2
7x = -13
x = -13/7